OpenStudy (anonymous):

Which value for x makes the sentence true? 8x - 4 = 2x + 20 0 6 8 4

OpenStudy (mathmath333):

\(\Large\tt \begin{align} \color{black}{8x - 4 = 2x + 20\\ 8x-2x = 20+4\\ 6x=24\\ x=\dfrac{24}{6}\\ x=4 }\end{align}\)